# PerturBench: Benchmarking Machine Learning Models for Cellular Perturbation Analysis

### v1

@@ Abstract
We present a comprehensive framework for predicting the effects of perturbations in single cells, designed to standardize benchmarking in this rapidly evolving field. Our framework, PerturBench, includes a user-friendly platform, diverse datasets, metrics for fair model comparison, and detailed performance analysis. Extensive evaluations of published and baseline models reveal limitations like mode or posterior collapse, and underscore the importance of rank metrics that assess the ordering of perturbations alongside traditional measures like RMSE. Our findings show that simple models can outperform more complex approaches. This benchmarking exercise sets new standards for model evaluation, supports robust model development, and advances the potential of these models to use high-throughput and high-content genetic and chemical screens for disease target discovery.
@@ 1 Introduction
Experiments where biological systems such as cell lines are perturbed via chemical treatment or genetic modification can help unravel the causal drivers of disease and identify promising therapeutics. Advances in CRISPR technology and automation have enabled these experiments, which we refer to as perturbation screens, to be conducted at scale with thousands to hundreds of thousands of perturbations applied in parallel in a single experiment (Shalem et al.,, 2014). These perturbation screens have been coupled with modern sequencing technology (see e.g. Lowe et al.,, 2017, for an overview) that can measure RNA expression profiles at single cell resolution, creating a rich understanding of perturbation effects (Tang et al.,, 2009; Macosko et al.,, 2015; Adamson et al.,, 2016; Dixit et al.,, 2016; Bock et al.,, 2022; Srivatsan et al.,, 2020). However, measuring these perturbation effects for the tens of thousands of currently known genes or the $10^{60}$ known drug-like chemicals across different tissues and cell types is prohibitively expensive with current technologies, especially for combinations of perturbations (Replogle et al.,, 2022; Reymond,, 2015). Thus, there is an increased interest in computational approaches that can predict the effects of perturbations on gene expression.
@@
Specifically, researchers have developed models that can generate counterfactual, out of sample (oos) predictions of perturbation effects (Gavriilidis et al.,, 2024). One use case, which we call covariate transfer, involves training a model on perturbation effects measured in a set of covariates (i.e. cell lines) and predicting those effects in another covariate where they were never observed. Another use case, which we call combo prediction, involves training a model on individual perturbation effects and predicting the effects of multiple perturbations in combination. The ultimate goal of these models is to enable in-silico screening across the vast space of unobserved experimental perturbations and pave the way for targeted disease treatment strategies.
@@
It is difficult to judge the performance of published perturbation effect prediction models against each other due to differences in benchmarking datasets and metrics. The NeurIPS 2023 perturbation prediction challenge helped address this issue by providing a chemical perturbation dataset with a model evaluation framework (Burkhardt et al.,, 2023). However, the challenge was limited to a single covariate transfer dataset/task and models were evaluated using a single metric that may not fully capture performance. Additionally, the challenge did not benchmark published models. The sc-perturb database provides datasets with unified metadata, but does not attempt to benchmark models (Peidli et al.,, 2024).
@@
In this work, we provide a rigorous quantitative assessment of the field of perturbation response modelling by: (1) introducing a highly modular and user-friendly framework in the form of a code base [^1] Available at: https://github.com/altoslabs/perturbench for model development and evaluation, (2) curating diverse perturbational datasets and defining biologically relevant tasks, (3) defining a set of metrics that enable comparison of different models on an equal basis, (4) performing extensive evaluation of existing perturbation models and their individual components across different datasets. The evaluation took more than 4 thousands hours of modern servers with GPUs.
@@
We reproduce key components of published models that cover a spectrum of architectures, and evaluate them against each other and strong baseline models. We specifically test the models on difficult tasks, simulating how they will be deployed in real-life contexts. Our findings reveal that some widely used models are prone to “mode” or “posterior” collapse (see Appendix D.1 for more details). Since a common use-case of these models is to run in-silico screens that rank perturbations by a desired effect (i.e. reversing a disease state) (Van de Sande et al.,, 2023), we propose rank metrics complementary to traditional measures of model fit (e.g. root mean squared error (RMSE)) that specifically capture model collapse. In addition, we demonstrate that models with simple architectures can outperform some of the more sophisticated models.
@@ 2 Datasets and Tasks
Many published models have been evaluated on relatively small datasets, with a split where most of the data is used for training. However, in a real-world setting, we often have complex datasets that only contain a small fraction of the perturbation effects we are interested in predicting. Thus, we select datasets and tasks that mirror these real-world challenges. We select three published datasets, Norman et al., (2019), Srivatsan et al., (2020), and McFaline-Figueroa et al., (2024), which include at least 100 perturbations and cover a diversity of perturbation modalities (chemical vs genetic), combinatorial perturbations, and dataset sizes. We provide a cursory overview in Table 1 and more information in Appendix E.1. Here, we define a biological state as a unique set of covariates that we plan to model (i.e. cell line). Details of the preprocessing of the datasets are in the Appendix E.3.
@@
Norman19 287 131 genetic 1 91,168 combo prediction
We create a covariate transfer split for the Srivatsan20 dataset and a combo prediction task for the Norman19 dataset. In addition, we study two scenarios: data scaling and imbalanced data. In the former, we benchmark model performance with increasing number of additional data. In the later, we simulate a situation where data of some covariate type is more abundant than that of another. Details of the experiments are in sections 5.3 and 5.4 respectively. The aim of both scenarios is to simulate how models will be deployed in practice, where there are often complex covariates, imbalanced datasets, and/or large amounts of missing data (see e.g. Edwards,, 2024). Additional details about data splitting implementation can be found in the Appendix E.4.
@@ 3.1 Modeling counterfactuals
Matching treated outcomes to controls is a common approach to identify treatment effects (see e.g. Stuart,, 2010, Section 1.3 for a historical summary). In the context of perturbation effect prediction, matching of perturbed cells with controls has been used by a variety of published models such as GEARS (Roohani et al.,, 2023), scGPT (Cui et al.,, 2024), and scFoundation (Hao et al.,, 2023). However, the effectiveness and validity of matched controls depend on fulfilling certain assumptions to ensure measuring a causal effect. One such assumption is ignorability, which posits that once covariates are controlled for, no residual confounding effects should influence the comparison between control and treatment groups (see e.g. Stuart,, 2010). This can include, for instance, ensuring that the control cell is from the same cell type, experiment or batch. Some more complex methods involve using optimal transport to identify the control cell most likely to transition into a given perturbed cell (Jiang et al.,, 2024) or even to use optimal transport to assist in perturbation prediction (Bunne et al.,, 2023). However, due to limited scalability of common algorithms to solve the optimal transport problem, these models have not been applied to larger datasets with more than 100 perturbations.
@@ 3.2 Models for benchmarking
In this paper we implement a range of perturbation response models inspired by some of the state-of-the-art models in the field: CPA, Biolord and SAMS-VAE. Our aim is to assess the core modeling component behind each model, such as the adversarial classifier for CPA, and thus there are differences between some of our implementations and the published versions, see Appendix E.5. We therefore refer to our implementations with a following the model name (i.e. CPA). We further encode gene expression values using scGPT, a single cell gene expression foundation model (Cui et al.,, 2024), and used the resulting embeddings as inputs to CPA and our latent additive baseline model. Our aim is to gauge whether using embeddings from a pretrained foundation model would improve performance. As many prior studies lacked strong baseline models, we also implement and benchmark the following baselines.
@@ 4 Benchmarking
Existing perturbation response modeling studies have used different metrics for different models, making comparison between models difficult. We develop a standardized benchmarking approach with a suite of metrics together with a simple API to facilitate evaluation and comparability of established and new perturbation models, see Appendix B for details.
@@ 4.2 Metrics
We select metrics that capture the performance of perturbation response prediction models in real world applications, specifically using RMSE (as recommended by Ji et al., 2023) to compare the average predicted cell states to the observed cell states. We also use cosine similarity to assess the similarity of the predicted versus observed perturbation effects as measured by LogFCs. However, as we show in Appendix D.1, these “global” fit-based metrics do not fully capture all aspects of a model’s performance. Hence, we introduce a set of rank-based metrics that can be seen as measures of precision.
@@ 4.2.1 Rank Metrics
We find it difficult to use existing infromation retrieval metrics such as the mean reciprocal rank (Radev et al.,, 2002) or mean average precision because we do not know in advance what the desired gene expression state and thus cannot partition perturbations into relevant and not relevant classes for these retrieval metrics. Hence, we introduce a rank-based metric that measures how well the predicted perturbations are ordered. Specifically, for a given observed perturbation, the prediction for that perturbation should be more similar than predictions for other perturbations. The rank metric is computed on a per perturbation basis:
@@ 4.3 Benchmarking Rules
The performance of each model critically depends on the judicious choice of hyperparameters. To identify suitable hyperparameters for each model, we run hyperparameter optimization (HPO) for each model on each dataset, task and scenario. Specifically, we use optuna (Akiba et al.,, 2019) with the default tree-structured Parzen estimator (Bergstra et al.,, 2011) to sample hyperparameters. Each model is run on a node with at least 64GB of RAM and one Nvidia A10 GPU; for the biggest dataset and models even stronger nodes are used. For every model we carry out training with at least 60 hyperparameter configurations and we describe the specific hyperparameters and their ranges in the Appendix E.6. For the best configuration we run model training four additional times with different seeds to assess stability of training. Error bars represent the standard deviation of model performance. To balance the two objectives, RMSE and corresponding ranking, we carried out a pilot HPO run and found that using a combination of the RMSE and 0.10.1$\cdot$0.1 ⋅ rank results in good overall performance. We include pilot run details in Appendix E.6. Hence, for HPO we define the loss function $\mathcal{L}_{\textsc{HPO}}=\text{RMSE}+0.1\cdot\mathrm{rank}_{\mathrm{RMSE}}$.
@@ 5.1 Predicting Drug Effects Across Cell Types
The results are summarized in Table 2. Interestingly, we see that the baseline models tend to outperform more sophisticated models. The latent additive model (LA) with scGPT embeddings performs strongest overall, indicating that scGPT embeddings provide rich representations for predicting perturbation effects. This is followed by the default latent additive model.
The decoder-only baseline (Decoder) that uses both perturbations and covariates does not achieve the same cosine LogFC but shows similar performance in terms of rank metrics. Both models outperform the more sophisticated CPA, SAMS-VAE and BioLord models, both in terms of the standard cosine and RMSE metrics and the ranking metrics. This suggests that at least in this limited setting, simple models relying on control matching approaches perform better than models that contain sophisticated disentanglement architectures. We investigate disentanglement in CPA by removing the adversarial classifier from CPA (CPA noAdv), which increased the variability in the benchmarking metrics and seemed to slightly improve on the average performance. This is a surprising observation and suggests that CPA’s perturbation encoder is able to learn a meaningful representation even without the adversarial classifier.
It is worth noting that the decoder-only baseline that only uses covariates and has no perturbation information, (Decoder (Cov)) achieves a fairly high cosine similarity and low RMSE. This suggests that it is possible for a model to find a single expression vector that is similar to all perturbations in a given cell type, which again highlights the need for rank metrics that assess whether models can correctly order perturbations. See Appendix D.1 for a more detailed assessment.
CPA performs worse than the linear (Linear) baseline model in the rank metrics, suggesting that CPA does not predict the unique effects of each perturbation as well. SAMS-VAE underperforms on all metrics, with rank metrics close to 0.5 suggesting it is not able to order perturbations better than random. Biolord shows high variance between training runs, suggesting the model initialization plays a large role in model performance for this task.
@@
SAMS-VAE $0.13\pm 1\times 10^{-3}$ $0.060\pm 1\times 10^{-3}$ $0.49\pm 3\times 10^{-2}$ $0.45\pm 2\times 10^{-2}$

### v2

@@ Abstract
We present a comprehensive framework for predicting the effects of perturbations on cellular states, designed to standardize benchmarking in this rapidly evolving field. Our framework, PerturBench, includes a user-friendly platform, diverse datasets, metrics for fair model comparison, and detailed performance analysis. Extensive evaluations of published and baseline models reveal limitations like mode or posterior collapse, and underscore the importance of rank metrics that assess the ordering of perturbations alongside traditional measures like RMSE. Our findings show that simple models can outperform more complex approaches. This benchmarking exercise sets new standards for model evaluation, supports robust model development, and advances the potential of these models to use high-throughput and high-content genetic and chemical screens for disease target discovery.
@@ 1 Introduction
Experiments in which biological systems, such as cell lines, are perturbed through chemical treatments or genetic modifications, can help unravel the causal drivers of diseases and identify promising therapeutics. Advances in CRISPR technology and automation have enabled these experiments, which we refer to as perturbation screens, to be conducted at scale with up to hundreds of thousands of perturbations applied in parallel in a single experiment (Shalem et al.,, 2014). These perturbation screens have been coupled with modern RNA-sequencing technology (see e.g. Lowe et al.,, 2017, for an overview) that can measure gene expression profiles at single cell resolution, creating a rich understanding of perturbation effects (Tang et al.,, 2009; Macosko et al.,, 2015; Adamson et al.,, 2016; Dixit et al.,, 2016; Bock et al.,, 2022; Srivatsan et al.,, 2020). However, measuring the effect of perturbing the tens of thousands of currently known genes or the $10^{60}$ known drug-like chemicals across different tissues and cell types is prohibitively expensive with current technologies, especially for combinations of perturbations (Replogle et al.,, 2022; Reymond,, 2015). Thus, there is an increased interest in computational approaches that can predict the effects of perturbations on gene expression.
@@
Specifically, researchers have developed models that can generate counterfactual, out of sample (oos) predictions of perturbation effects (Gavriilidis et al.,, 2024). One use case, which we call covariate transfer, involves training a model on perturbation effects measured in a set of covariates (i.e. cell lines) and predicting those effects in another covariate where they were never observed. Another use case, which we call combo prediction, involves training a model on individual perturbation effects and predicting the effects of multiple perturbations in combination. The ultimate goal of these models is to enable in-silico screening across the vast space of unobserved perturbations and pave the way for targeted disease treatment strategies.
@@
It is difficult to judge the performance of published perturbation effect prediction models against each other due to differences in benchmarking datasets and metrics. The NeurIPS 2023 perturbation prediction challenge helped address this issue by providing a chemical perturbation dataset with a model evaluation framework (Burkhardt et al.,, 2023). However, the challenge was limited to a single dataset on the covariate transfer task and models were evaluated using a single metric that may not fully capture performance. Additionally, the challenge did not benchmark published models. The sc-perturb database provides datasets with unified metadata, but does not attempt to benchmark models (Peidli et al.,, 2024). Ahlmann-Eltze et al., (2024) and Wenteler et al., (2024) assessed the performance of single cell foundation models and GEARS on only the combo prediction and unseen perturbation prediction tasks, and both studies used MSE as their evaluation metric, which again may not capture key aspects of model performance.
@@
In this work, we provide a rigorous quantitative assessment of the field of perturbation response modelling by: (1) introducing a highly modular and user-friendly framework in the form of a code base for model development and evaluation, (2) curating diverse perturbational datasets and defining biologically relevant tasks, (3) defining a set of metrics that enable comparison of different models on an equal basis, (4) performing extensive evaluation of existing perturbation models and their individual components across different datasets. Figure 1 illustrates our approach.
Figure 1: A) Single cell perturbational datasets at multiple scales. B) Biologically relevant covariate transfer and combinatorial prediction data splits. C) Dataloaders support two training modes: 1) supervised mode which involves mapping a sampled control cell to a perturbed cell and 2) unsupervised (or autoencoder) mode which maps a perturbed cell to itself. D) A model zoo with modular components such as relevant baseline models, adversarial loss, perturbation sparsity, and others. E) Standardized benchmarking suite supporting flexible pipelines and metrics for evaluating models
@@
We reproduce key components of published models that cover a spectrum of architectures, and evaluate them against each other and strong baseline models. We specifically test the models on difficult tasks, simulating how they will be deployed in real-life contexts. Our findings reveal that some widely used models are prone to “mode” or “posterior” collapse (see Appendix C.1 for more details). Since a common use-case of these models is to run in-silico screens that rank perturbations by a desired effect (i.e. reversing a disease state) (Van de Sande et al.,, 2023), we propose rank metrics complementary to traditional measures of model fit (e.g. root mean squared error (RMSE)) that specifically capture model collapse. In addition, we demonstrate that models with simple architectures can outperform some of the more sophisticated models. In total, our benchmarking required more than 4 thousands GPU hours.
@@ 2 Datasets and Tasks
Many published models have been evaluated on relatively small datasets, with a split where most of the data is used for training. However, in a real-world setting, we often have complex datasets that only contain a small fraction of the perturbation effects we are interested in predicting. Thus, we select datasets and tasks that mirror these real-world challenges. We select five published datasets, Norman et al., (2019), Srivatsan et al., (2020), Frangieh et al., (2021), McFaline-Figueroa et al., (2024), and Jiang et al., 2024a which include at least 100 perturbations and cover a diversity of perturbation modalities (chemical vs genetic), combinatorial perturbations, dataset sizes, and covariates. We provide a cursory overview in Table 1 and more information in Appendix D.1. Here, we define a biological state as a unique set of covariates that we plan to model (i.e. cell line). Details of the preprocessing of the datasets are in the Appendix D.3.
@@
Frangieh21 248 0 genetic 3 218,331 covariate transfer
Jiang24 219 0 genetic 30 1,628,476 covariate transfer
@@
Norman19 155 131 genetic 1 91,168 combo prediction
We create a covariate transfer split for the Srivatsan20, Frangieh21 and Jiang24 datasets as well as a combo prediction task for the Norman19 dataset. In addition, we study two scenarios: data scaling and imbalanced data. In the former, we benchmark model performance with increasing number of additional data. In the later, we simulate a situation where data of some covariate type is more abundant than that of another. Details of the experiments are in sections 5.4 and 5.5 respectively. The aim of both scenarios is to simulate how models will be deployed in practice, where there are often complex covariates, imbalanced datasets, and/or large amounts of missing data (see e.g. Edwards,, 2024). Additional details about data splitting implementation can be found in the Appendix D.4.
@@ 3.1 Modeling counterfactuals
Matching treated outcomes to controls is a common approach to identify treatment effects (see e.g. Stuart,, 2010, Section 1.3 for a historical summary). In the context of perturbation effect prediction, matching of perturbed cells with controls has been used by a variety of published models such as GEARS (Roohani et al.,, 2023), scGPT (Cui et al.,, 2024), and scFoundation (Hao et al.,, 2023). However, the effectiveness and validity of matched controls depend on fulfilling certain assumptions to ensure measuring a causal effect. One such assumption is ignorability, which posits that once covariates are controlled for, no residual confounding effects should influence the comparison between control and treatment groups (see e.g. Stuart,, 2010). This can include, for instance, ensuring that the control cell is from the same cell type, experiment or batch. Some more complex methods involve using optimal transport to identify the control cell most likely to transition into a given perturbed cell (Jiang et al., 2024b, ) or even to use optimal transport to assist in perturbation prediction (Bunne et al.,, 2023). However, due to limited scalability of common algorithms to solve the optimal transport problem, these models have not been applied to larger datasets with more than 100 perturbations.
@@ 3.2 Models for benchmarking
In this paper we implement a range of perturbation response models inspired by some of the state-of-the-art models in the field: CPA, Biolord, SAMS-VAE and GEARS. Our aim is to assess the core modeling component behind each model, such as the adversarial classifier for CPA, and thus there are differences between some of our implementations and the published versions, see Appendix D.5. We therefore refer to our implementations with a following the model name (i.e. CPA). We further encode gene expression values using scGPT, a single cell gene expression foundation model (Cui et al.,, 2024), and used the resulting embeddings as inputs to CPA and our latent additive baseline model. Our aim is to gauge whether using embeddings from a pretrained foundation model would improve performance. As many prior studies lacked strong baseline models, we also implement and benchmark the following baselines.
@@ 4 Benchmarking
Existing perturbation response modeling studies have used different metrics for different models, making comparison between models difficult. We develop a standardized benchmarking approach with a suite of metrics together with a simple API to facilitate evaluation and comparability of established and new perturbation models, see Appendix A for details.
@@ 4.2 Metrics
We select metrics that capture the performance of perturbation response prediction models in real world applications, specifically using RMSE (as recommended by Ji et al., 2023) to compare the average predicted cell states to the observed cell states. We also use cosine similarity to assess the similarity of the predicted versus observed perturbation effects as measured by LogFCs. However, as we show in Appendix C.1, these “global” fit-based metrics do not fully capture all aspects of a model’s performance. Hence, we introduce a set of rank-based metrics that can be seen as measures of precision.
@@ 4.2.1 Rank Metrics
We find it difficult to use existing information retrieval metrics such as the mean reciprocal rank (Radev et al.,, 2002) or mean average precision because we do not know in advance what the desired gene expression state and thus cannot partition perturbations into relevant and not relevant classes for these retrieval metrics. Hence, we introduce a rank-based metric that measures how well the predicted perturbations are ordered. Specifically, for a given observed perturbation, the prediction for that perturbation should be more similar than predictions for other perturbations. The rank metric is computed on a per perturbation basis:
@@ 4.3 Benchmarking Rules
The performance of each model critically depends on the judicious choice of hyperparameters. To identify suitable hyperparameters for each model, we run hyperparameter optimization (HPO) for each model on each dataset, task and scenario. Specifically, we use optuna (Akiba et al.,, 2019) with the default tree-structured Parzen estimator (Bergstra et al.,, 2011) to sample hyperparameters. Each model is run on a node with at least 64GB of RAM and one Nvidia A10 GPU; for the biggest dataset and models even stronger nodes are used. For every model we carry out training with at least 60 hyperparameter configurations and we describe the specific hyperparameters and their ranges in the Appendix D.6. For the best configuration we run model training four additional times with different seeds to assess stability of training. Error bars represent the standard deviation of model performance. To balance the two objectives, RMSE and corresponding ranking, we carried out a pilot HPO run and found that using a combination of the RMSE and 0.10.1$\cdot$0.1 ⋅ rank results in good overall performance. We include pilot run details in Appendix D.6. Hence, for HPO we define the loss function $\mathcal{L}_{\textsc{HPO}}=\text{RMSE}+0.1\cdot\mathrm{rank}_{\mathrm{RMSE}}$.
@@ 5.1 Predicting Drug Effects Across Cell Types
The results are summarized in Table 2. Interestingly, we see that the baseline models tend to outperform more sophisticated models. The latent additive model (LA) with scGPT embeddings performs strongest overall, indicating that scGPT cell embeddings provide rich representations for predicting perturbation effects. This is followed by the default latent additive model. SAMS-VAE is the second best performing model, performing similar to or better than other published models on all metrics.
The decoder-only baseline (Decoder) that uses both perturbations and covariates does not achieve the same cosine LogFC but shows similar performance in terms of rank metrics. Both models outperform the more sophisticated CPA, and BioLord models, both in terms of the standard cosine and RMSE metrics and the ranking metrics. We investigate disentanglement in CPA by removing the adversarial classifier from CPA (CPA noAdv), which increased the variability in the benchmarking metrics and seemed to slightly improve on the average performance. This is a surprising observation and suggests that CPA’s perturbation encoder is able to learn a meaningful representation even without the adversarial classifier.
It is worth noting that the decoder-only baseline that only uses covariates and has no perturbation information, (Decoder (Cov)) achieves a fairly high cosine similarity and low RMSE. This suggests that it is possible for a model to find a single expression vector that is similar to all perturbations in a given cell type, which again highlights the need for rank metrics that assess whether models can correctly order perturbations. See Appendix C.1 for a more detailed assessment.
CPA performs worse than the linear (Linear) baseline model in the rank metrics, suggesting that CPA does not predict the unique effects of each perturbation as well. Biolord shows high variance between training runs, suggesting the model initialization plays a large role in model performance for this task.
@@
SAMS-VAE $0.44\pm 1\times 3^{-3}$ $0.023\pm 8\times 10^{-5}$ $0.17\pm 1\times 10^{-2}$ $0.17\pm 1\times 10^{-2}$
@@
5.2 Generalizing from less complex to more complex biological systems
[…]
The results are summarized in Table 3. The latent additive model, CPA both with and without scGPT embeddings, and SAMS-VAE are best at predicting the expression similarity as measured by the cosine LogFC and RMSE. However, CPA and SAMS-VAE do not perform well on either rank metric, while the latent additive model and BioLord perform similarly well on the rank metric. The decoder only model performs best on both rank metrics while underperforming CPA and the latent additive model on the cosine logFC and RMSE metrics.
Table 3: Results of a covariate transfer experiment (mean $\pm$ one standard deviation) generalizing from less complex biological systems to a more complex co-culture system in the Frangieh21 dataset. Results are reported as mean $\pm$ one standard deviation and best performance per metric is indicated in bold.
Model Cosine RMSE Cosine RMSE
LogFC mean LogFC rank mean rank
CPA $0.17\pm 1\times 10^{-2}$ $0.027\pm 6\times 10^{-4}$ $0.41\pm 2\times 10^{-2}$ $0.42\pm 1\times 10^{-2}$
CPA (scGPT) $0.16\pm 1\times 10^{-2}$ $0.026\pm 2\times 10^{-3}$ $0.46\pm 7\times 10^{-3}$ $0.48\pm 3\times 10^{-2}$
SAMS-VAE $0.15\pm 2\times 10^{-2}$ $0.026\pm 2\times 10^{-4}$ $0.48\pm 2\times 10^{-2}$ $0.46\pm 2\times 10^{-2}$
BioLord $0.12\pm 9\times 10^{-3}$ $0.027\pm 2\times 10^{-4}$ $0.29\pm 3\times 10^{-2}$ $0.21\pm 4\times 10^{-2}$
LA $\mathbf{0.17\pm 6\times 10^{-3}}$ $\mathbf{0.024\pm 4\times 10^{-4}}$ $0.26\pm 1\times 10^{-2}$ $0.21\pm 2\times 10^{-2}$
LA (scGPT) $\mathbf{0.18\pm 6\times 10^{-3}}$ $\mathbf{0.024\pm 6\times 10^{-5}}$ $0.27\pm 1\times 10^{-2}$ $0.24\pm 1\times 10^{-2}$
Decoder $0.10\pm 2\times 10^{-3}$ $\mathbf{0.025\pm 4\times 10^{-5}}$ $\mathbf{0.21\pm 5\times 10^{-3}}$ $\mathbf{0.15\pm 4\times 10^{-4}}$
Linear $0.0081\pm 4\times 10^{-4}$ $0.043\pm 7\times 10^{-5}$ $0.24\pm 9\times 10^{-4}$ $0.30\pm 2\times 10^{-3}$